OpenStudy (anonymous):

a positive two digit number is represented by 10t+u. the sum of the squares of the digits is 26. if the number is decreased by the number with its digits reversed, the result if 36. find the two digit number...

OpenStudy (anonymous):

answer choices: 51, 26, 38, 67

OpenStudy (anonymous):

t^2 + u^2 = 26 10t + u - 10u - t = 9t - 9u = 36 9 (t-u) = 36 t - u = 4

OpenStudy (anonymous):

t = u + 4 (u + 4)^2 + u^2 = 26 u^2 + 8u + 16 + u^2 = 26 2u^2 + 8u -10 = 0 u^2 + 4u - 5 = 0 (u-1) (u + 5) = 0 u = 1, -5

OpenStudy (anonymous):

u must be 1 since it's a digit. t^2 + 1 = 26 t^2 = 25 t = -5, 5 t must be 5 since it's a digit. answer is 51.

OpenStudy (anonymous):

Do you understand, ma'am?

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

call me emily.. ma'am makes me feel really old :P

OpenStudy (anonymous):

ok emily do you understand?

OpenStudy (anonymous):

yes this really helped.. thanks again..